Beagle Beauties engages in the development, manufacture, and sale of a line of cosmetics designed to make your dog look glamorous. Below you will find selected information necessary to compute some valuation estimates for the firm. Assume the values provided are from year-end 2015. Also assume that the firm’s equity beta is 1.40, the risk-free rate is 2.40 percent, and the market risk premium is 6.8 percent.

Dividends per share $ 3.12
Return on equity 12.00 %
Book value per share $ 17.50
Earnings Cash Flow Sales
2015 value per share $ 5.90 $ 7.05 $ 26.10
Average price multiple 14.00 9.69 2.72
Forecasted growth rate 13.66 % 11.86 % 7.88 %



The required return is 11.92 percent. Use the clean surplus relationship to calculate the share price for Beagle Beauties with the residual income model.
Answer #1

Using residual Income model,

Residual Income Per share in Year 1 = Earnings Per Share - Required return on equity * Beginning BVPS = 5.9- 12%*17.5

= $ 3.8

Cost of Equity , r = Rf+ beta * MRP = 2.4% + 1.4*6.8%= 11.92 %

Residual Income Valuation, V = B + [(ROE-r) /(r-g) * B ]= [17.5 + 17.5*(12-11.92)/(11.92-11.86)] = $ 40.833
Answer #2


Clean surplus means the dividends are not included Clean surplus relationship is as follows: Ending book value Beginning book